OpenStudy (anonymous):

solve using the multiplication principle. give the answer in set-builder notation. -4/5

OpenStudy (shadowfiend):

We have: \[-\frac{4}{5} \leq -6x \] We can divide by -6 on either side, but remember that dividing by a negative flips the inequality; so we have: \[-\frac{4}{-5\cdot 6} \geq x\] \[-\frac{4}{30} \geq x\] Can you simplify that fraction and write it the inequality set-builder notation?

OpenStudy (anonymous):

So shadowfend, would the answer be x > greater than or equal to -2/15

OpenStudy (shadowfiend):

Not quite, I made a sign error up there. The negatives cancel. And you said greather than or equal to whereas that should be less than or equal to. So: \[x \leq \frac{2}{15}\]

OpenStudy (shadowfiend):

Do you know how to put that in set-builder notation?

OpenStudy (anonymous):

no please help me

OpenStudy (anonymous):

help please translate to an inequality a number is at least 5 use x as the variable

OpenStudy (anonymous):

translate to an inequality use the variable x the number of people in the chess club is less than or equal to 19

OpenStudy (anonymous):

translat to an inequality the average speed, s, was between 130 and 140 mph

OpenStudy (anonymous):

translate to an inequality use the variable x the price of a baseball glove is at most $38.95

OpenStudy (anonymous):

corretta pls help translate to an inequality use the variable x the price of a baseball glove is at most $38.957 minutes ago

OpenStudy (anonymous):

loco85 pls help translat to an inequality the average speed, s, was between 130 and 140 mph

OpenStudy (anonymous):

gothic warrior pls help help please translate to an inequality a number is at least 5 use x as the variable

OpenStudy (anonymous):

quantum torch pls help translate to an inequality use the variable x the price of an amusement park ticket is greater than or equal to $18 the inequality is blank____ blank_____ $____ simplify your answer put the variable on the left side of the inequality